# Network-Level Prompt and Trait Leakage in Local Research Agents

### v2

@@ C.3 Inference Configuration
We evaluate our inference attacks using LLMs as an inference engine for few-shot ICL. A summary of ICL configurations is provided in Table 12, and the full ICL prompt templates are in Appendix G and Appendix I.
Prompt Recovery. To recover the user’s original prompt from a trace, we use a 5-shot ICL configuration. Each ICL prompt consists of five (trace, prompt) pairs, where the trace includes an ordered domain sequence and associated timing information, followed by an unseen trace for prediction. The format is fully natural language, with traces presented as newline-separated domain lists (and timing metadata when present). Examples are drawn from distinct topics to prevent overfitting and to test generalization across different intent types. This format encourages the model to infer high-level task semantics from the structure, content, and temporal progression of visited domains.
@@
| Task | ICL Shots | Output Format | Prompt Details |
| Prompt Recovery | 5 | Natural language query | Domain trace prompt |
| Trait Inference | 3 | Structured list (-Trait: Value) | Domain trace trait list |
Appendix D Additional Prompt Recovery Results
Effect of ICL Contrastive Configurations. Gao et al. [22] introduced contrastive ICL, showing that adding negative examples to the in-context setup can improve performance, even when negatives must be synthesized by the LLM. Since our datasets lack native negatives, we follow their procedure to generate them automatically.
In our baseline, each 5-shot prompt includes five (trace, prompt) pairs followed by a new trace for prediction (Figure 13). Contrastive variants augment these with one or three negatives, either in a simple setup or in our proposed Quality-Filtered (QF) contrastive ICL, where negatives are regenerated until their SBERT similarity to the original prompt falls below a threshold, ensuring sufficient dissimilarity.
@@ Appendix D Additional Prompt Recovery Results
Effect of Example Selection. We compare random versus embedding-based selection of ICL examples. In the random strategy, examples are sampled from the training set for each test instance. For embedding-based selection, all training and test traces are encoded with the all-MiniLM-L6-v2444Available via Hugging Face: https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2. Fine-tuned on 1B sentence pairs with a contrastive objective, this model maps each trace into a 384-dimensional dense vector space. sentence transformer, and the training examples with the highest cosine similarity to the test trace are chosen. As shown in Table 14, embedding-based selection does not improve prompt recovery accuracy and, in some metrics, performs slightly worse, suggesting that random sampling, by providing more diverse examples, is the more effective strategy.
@@
Appendix E Additional Virtual Persona Defense Results
Figure 8 and Figure 9 provide additional experimental results of the proposed virtual persona defense beyond the median similarity score reductions reported in the main text.
Appendix F OBELS Prompt Templates
Figure 10 shows the prompt used in the OBELS framework to convert natural language queries into structured semantic triplets of the form (intent, source_type, entity). These triplets capture the user’s goal, the type of resource sought, and the core entity involved. Prompts are drawn from the TREC Session Track 2014 dataset [13] and span a variety of user intents, including planning, learning, comparing, and exploring. For brevity, some prompts are abbreviated with ellipses (“…”) to fit within the template box, though the actual prompts used were complete.
Figure 11 then shows the complementary OBELS scoring prompt, which evaluates behavioral similarity between reconstructed and original prompts. The model is instructed to align triplets and score their correspondence along four dimensions: functional equivalence, domain type, semantic alignment, and entity granularity. This structured format ensures consistent evaluation across diverse prompt re.
Appendix G 5-Shot ICL Prompt Templates (Prompt Recovery)
Figure 13 presents the 5-shot ICL prompt used for reconstructing user intent from observed domain traces. Each example contains a list of domains followed by the original prompt that likely triggered them. The model is prompted to generate a plausible user prompt consistent with the browsing behavior.
Appendix H Prompt Template for Generating Persona-Based Queries
Figure 15 shows the full prompt used to elicit naturalistic web search queries from a simulated user. The goal is to mimic realistic multi-session browsing behavior based on a given persona, with selected traits subtly reflected in the generated queries.
Appendix I 3-Shot ICL Prompt Templates (Trait Inference)
Figure 16 illustrates the 3-shot ICL prompt used for inferring user traits from browsing data. Each example pairs a list of visited domains with a set of traits across 32 predefined dimensions. The model is instructed to extract as many traits as possible based solely on behavioral cues.
The examples used in the few-shot template are based on persona_000, persona_003, and persona_004. For brevity, we omit some traits from persona_003 in Figure 16. In the actual prompt used during inference, the full list of traits for persona_003 is included.

### v3

@@ C.3 Inference Configuration
We evaluate our inference attacks using LLMs as an inference engine for few-shot ICL. A summary of ICL configurations is provided in Table 12.
Prompt Recovery. Each ICL prompt consists of five (trace, prompt) pairs, where the trace includes an ordered domain sequence and associated timing information, followed by an unseen trace for prediction. The format is fully natural language, with traces presented as newline-separated domain lists (and timing when present). Examples are drawn from distinct topics to prevent overfitting and to test generalization across different intent types. This format encourages the model to infer high-level task semantics from the structure, content, and temporal progression of visited domains.
@@
Appendix D Additional Prompt Recovery Results
Effect of ICL Configurations. Contrastive ICL[gao2024customizing] showed that adding negative examples to the in-context setup can improve performance, even when negatives must be synthesized by the LLM. In our baseline, each 5-shot prompt includes five (trace, prompt) pairs followed by a new trace for prediction. Contrastive variants augment these with one or three negatives, either in a simple setup or in our proposed Quality-Filtered (QF) contrastive ICL, where negatives are regenerated until their SBERT similarity to the original prompt falls below a threshold, ensuring sufficient dissimilarity.
Effect of Example Selection. We compare random versus embedding-based selection of ICL examples. In the random strategy, examples are sampled from the training set for each test instance. For embedding-based selection, all training and test traces are encoded with the all-MiniLM-L6-v2111Fine-tuned on 1B sentence pairs with a contrastive objective, this model maps each trace into a 384-dimensional vector space. sentence transformer, and the training examples with the highest cosine similarity to the test trace are chosen. As shown in Table 14, embedding-based selection does not improve prompt recovery accuracy and, in some metrics, performs slightly worse, suggesting that random sampling, by providing more diverse examples, is the more effective strategy.